OpenStudy (anonymous):

Find the exact value of tan (arcsin (2/5)).

zepdrix (zepdrix):

OOoo fun problem :) I always enjoy these XD So let's think about it piece by piece hmmm

zepdrix (zepdrix):

\[\sin^{-1} (2/5)= \theta\] Is the same thing as writing\[\sin (\theta) = 2/5\] since sine and arcsine are inverse functions. does that part make sense? :D

OpenStudy (anonymous):

I never really understood the ponto of θ lol

zepdrix (zepdrix):

 Next thing we want to try to do, is to visualize what arcsin(2/5) looks like. So we draw a triangle!